Circular No. 6885

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1808.4–3658 = XTE J1808–369 P. Roche, University of Sussex; D. Chakrabarty, Massachusetts Insti-

tute of Technology (MIT); L. Morales-Rueda and R. Hynes, University of Sussex; S. M. Slivan, MIT; C. Simpson, Jet Propulsion Laboratory; and P. Hewett, Institute of Astronomy, Cambridge, report: “We have identified a possible optical counterpart for the millisecond x-ray pulsar SAX J1808.4–3658 (IAUC 6876, 6877, 6878). Observations with the 1-m Jacobus Kapteyn Telescope at La Palma and the 3.8-m U.K. Infrared Telescope at Mauna Kea on Apr. 18 UT detected a star that is not present to a limiting magnitude of V > 19 in the Digitized Sky Survey (DSS) image of the field (from a 1987 U.K. Schimdt plate). Its position is α = 18^{h}08^{m}27^{s}.5, δ = −36^{o}58′42′′ (equinox 2000.0), or 19′′ from the center of the 0′.6 (68-percent confidence) x-ray error circle from SAX (in’t Zand et al. 1998, A.Ap. 331, L25). Its position is also consistent with the smaller 25′′ uncertainty from the x-ray timing data (Chakrabarty and Morgan 1998,

http://xxx.lanl.gov/abs/astro-ph/?9804248). No other stars in our

V -band images of the SAX error circle showed significant variability with respect to the DSS image, to a limiting magnitude of ∼ 17.5. The magnitudes of the candidate are V = 16.6, R = 16.1, I = 15.6, J = 15.0, H = 14.4, K = 13.8, with an uncertainty of 0.2 mag in VRI and < 0.1 mag in JHK. We note that a fainter star (V ∼ 18) that is located 5′′ south-southeast of the candidate position is visible on both the DSS image and our JHK images and so is not our candidate. Spectroscopy and further photometry are urgently required to check whether the candidate is an unrelated background variable, and to look for fading as the x-ray source declines, as well as a possible 2-hr orbital modulation in the intensity. Observations with the Rossi X-Ray Timing Explorer indicate that the x-ray source is declining on a timescale of 2 weeks.”

SUPERNOVA 1998S IN NGC 3877 CCD photometry by D. Hanzl, Brno, Czech Republic, with a 0.4-m

reflector (G = comparison star GSC 3452.1093, V = 9.87 ± 0.04, B − V = +0.9 ± 0.06; T = Tycho comparison stars): Apr. 8.869 UT, V = 12.77 (G); 11.892, V = 12.91 (G); 13.832, V = 13.02 ± 0.02 (G); 20.895, V = 13.32 (T); 25.873, V = 13.49 ± 0.01, B − V = +0.96 ± 0.04 (T); 26.928, V = 13.54 ± 0.03, B − V = +0.91 ± 0.06 (G).

1998 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green